OpenStudy (anonymous):

if x = 7+[\sqrt{40}\]. find the value of [\sqrt{x}\]+1/[\sqrt{x}\]. pls give me an answer with steps or hints.

OpenStudy (kropot72):

Are these right? \[x=7+\sqrt{40}\] \[\frac{\sqrt{x}+1}{\sqrt{x}}\]